OpenStudy (anonymous):

What is the factored form of the following expression? d2 – 14d + 45 A. (d + 9)(d + 5) B. (d + 9)(d – 5) C. (d – 9)(d + 5) D. (d – 9)(d – 5) A, B, C, or D? Please explain (:

OpenStudy (anonymous):

1) D

OpenStudy (anonymous):

--> d2 -14d +45 = d2 - 9d-5d + 45 = d(d -9) -5(d-9)

OpenStudy (anonymous):

-> d2 -14d +45 = d2 - 9d-5d + 45 = d(d -9) -5(d-9) (d-9) (d-5)

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

you are always wel come:-)